The equilibrium constant for the formation of Cu(CN)4^2- is 2.0 x 10^-30. Calculate the value of pCu2+, or -log[Cu2+], if we were to dissolve 2.76 g of CuCl2 in 1.000 L of a solution 0.848 M in NaCN. The addition of CuCl2 does not alter the volume (the final volume is still 1.000 L).
1
Expert's answer
2020-07-10T05:25:10-0400

CuCl2 +4NaCN = Na2[Cu(CN)4] + 2NaCl

Cu(CN)42- = Cu2+ + 4CN-

K = [Cu2+][ CN-]4 / [Cu(CN)42-] = 2×10-30

Molecular weight of CuCl2 is 134.45 g/mol

Number of mol CuCl2 is = 2.76/134.45 mol =0.0205 mol

Number of mol NaCN is = 0.848 × 1000 = 848 mol  

 [Cu(CN)42-] = 0.0205 mol and CM = 0.0205/1000 = 2.05 × 10-5

Cu(CN)42- = Cu2+ + 4CN-     [Cu2+ = S, CN- =S]

K = S ×4S4 = 4S5

K×[Cu(CN)42-] = 4S5

S =[(2×10-30) × (2.05 × 10-5)/4]1/5 = 1.00049× 10-7

p[Cu2+] or -log[Cu2+] = -log(1.00049× 10-7) = 6.999 (Answer)